OpenStudy (anonymous):

I'm stuck on this trig substitution problem. integral of sqrt(x^2-64)/x dx I got the theta down to the integral of 8tan^3(x). So I did that and ended up with: 8((cos^2(x)/2) + log(sin(x)) Am I on the right track? Where do I go from here?

OpenStudy (whpalmer4):

Hmm, I don't think you did something correctly. How about using \(x=8\sec u\) and \(dx = 8\tan(u) \sec(u)\,du\) as your substitution?

OpenStudy (whpalmer4):

@ganeshie8 can help you with this, I have to call it a night...

ganeshie8 (ganeshie8):

I'll be here :) but I see he already got what he needs exactly :)

ganeshie8 (ganeshie8):

\(\large \int \frac{\sqrt{x^2-64}}{x} dx\) \(\large \int \frac{\sqrt{(8\sec u)^2-64}}{8\sec u } (8 \sec u \tan u du)\) \(\large \int \sqrt{(8\sec u)^2-64}~~(\tan u du)\) \(\large \int 8\sqrt{(\sec u)^2-1}~~(\tan u du)\) \(\large \int 8\sqrt{(\tan u)^2}~~(\tan u du)\) \(\large \int 8(\tan u)~~(\tan u du)\) \(\large \int 8 (\tan u)^2 du\)

ganeshie8 (ganeshie8):

it simplifies to that right ?

OpenStudy (anonymous):

Yep, I see what I did now. I forgot to take the root of tan^2

ganeshie8 (ganeshie8):

:) see if u can manage the rest u will need to rewrite tan in terms of sec and take the integral

ganeshie8 (ganeshie8):

also dont forget to convert u's to x's... good luck !

OpenStudy (anonymous):

Thanks! Hopefully I won't mess it up somewhere down the line. I'm pretty new at trig substitution. Let me see what I come up with.

OpenStudy (anonymous):

So you go back to the x= 8sec(u) formula and say u= sec^-1(x/8) right? I ended up with tan(sec^-1(x/8))-8sec^-1(x/8)

ganeshie8 (ganeshie8):

technically, thats correct... but that doesnt look neat right ? so to make it look neat, u need to draw a triangle.

ganeshie8 (ganeshie8):

\(\large \int 8 (\tan u)^2 du\) \(\large 8\int (\sec u)^2 - 1 du\) \(\large 8\tan u - 8u + c \)

ganeshie8 (ganeshie8):

you got that after taking integral right ?

ganeshie8 (ganeshie8):

to convert u's to x's, you need to draw a triangle

OpenStudy (anonymous):

Yes I did.

OpenStudy (anonymous):

Would you use x as the hypotenuse and 8 as the adjacent side? That's what it looks like according to my other answer.

ganeshie8 (ganeshie8):

right ! \(x = 8 \sec u\)

ganeshie8 (ganeshie8):